COLLISIONAL DEPOLARIZATION OF ZEEMAN COHERENCES IN THE ^{133}Cs 6p^{2}P_{3/2} LEVEL WITH ARGON MOLECULES: TWO-PHOTON POLARIZATION SPECTROSCOPY

\underline{S. B. BAYRAM}^{a}, S. KIN, M. WELSH and R. MARHATTA, Miami University, Physics Department, Oxford, OH 45056.

An experimental investigation of the Cs 6s^{2}S_{1/2} → 6p^{2}P_{3/2} → 10s^{2}S_{1/2} double resonance two-photon polarization spectrum has been made. In the investigation, a linear polarization degree was measured to extract the alignment depolarization cross section in the cesium 6p^{2}P_{3/2} level due to collisions between ground-level argon and excited 6p^{2}P_{3/2} cesium atoms over the Zeeman coherences. The spectra show strong depolarization of the Cs 6p^{2}P_{3/2} level. The measured alignment depolarization cross section value of 186 (58)Å^{2} is in good agreement with experiment and theoretical predictions.

aSupport by Research Corporation under the grant number CC6119C is gratefully acknowledged